• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Huber

    Log entry time 15:54:03 on September 27,1999

    Entry number 27270

    keyword=Day Shift Summary

    Run 1866
    9:07 - End run.

    Run 1867
    9:08 - Start run. Beam intensity continuing to drop steadily with time. Now at
    10.6uA.
    10:08 - End run. Beam intensity now down to 9.2uA.

    Wait for ReCesiation and NF3 treatment of source.
    Matt Poelker thinks we may have a QE hole at the present spot, and we should move the beam a few spot diameters. He can tell us when beam comes back up.

    Rebooting TS1. Bob Michaels has installed new diagnostic software for the next
    time the DAQ hangs. Read his log entry (27261) for instructions.

    11:15 - Beam back. Matt expects that we should get about 10uA now. During
    the previous injector tuning, the spot should have moved less than its diameter on
    the wafer, and so should not have greatly affected the polarization. Nonetheless,
    it seems like a good idea to remeasure it anyway.

    Eugene is now performing a Moller polarimeter measurement. Beam flux helicity
    assymetry is a high 3E-3. Polarization is lower now, 62.4%.

    Run 1868 - junk run.

    Run 1869
    13:41 - Start run. We are only getting 16uA.
    14:46 - End run.

    Run 1870
    14:47 - Start run. Beam intensity slowly dropping. Now 14.9uA.
    15:41 - End run.

    D+6%rad setting now completed.

    Insert LH target.

    Run 1871
    15:44 - Start run.

    Michael Kuss reboots target IOC